# Web Artifact Attacks Disrupt Vision Language Models

### v1

@@ 3.3 Attack Success Evaluation
Given an input image $x$ and an artifact $a$ placed at location $l$, we define the attack-modified image as $x^{a}=x\oplus a$, where $\oplus$ denotes artifact insertion. To verify that misclassification stems from the artifact itself, we introduce a masking-based validation step. We generate a masked version of the image, $x^{\neg l}$, where the region $l$ is occluded using a binary mask $M_{l}$.
@@
Figure 4: Success Rates of Artifact Attacks. Our artifact-based attacks are highly effective across different configurations: namely (a) Vision/Language Fusion, (b) Pretraining Dataset, and (c) Vision Encoder Architecture. Refer to Sec. 4.1 for further discussion..
@@ 4 Experiments
Datasets. We use CC12M [5] as our primary dataset for searching and extracting artifacts. Our extraction process yields 87k artifacts, but this approach is dataset-agnostic and can be applied to any large-scale image-text dataset, such as LAION [28, 29]. To evaluate the effectiveness of these artifacts, we conduct experiments across five distinct tasks spanning both human-related attributes and object/geography classification. For human-related tasks, we predict Gender (Male/Female) and Age (Child, Teen, Adult, Elderly) using the FairFace dataset [14], as well as Smiling using the CelebA dataset [19]. For non-human classification tasks, we predict aircraft models using the FGVC Aircraft dataset [20] and countries using the Country211 dataset [26]. For each downstream dataset, we use only 10 images per class to estimate the Artifact Success Rate (ASR), which as we demonstrate in Sec. 4.1, is sufficient to estimate ASR.
Metrics. We use the success rate as outlined in Sec. 3.3. In short, given a target class $c^{*}$, we measure attack effectiveness by evaluating whether the model misclassifies an image after the artifact that correlates with $c^{*}$ is introduced while ensuring that the image remains correctly classified when the artifact is masked. We also report the change of the confidence score in target class $c^{*}$ after pasting the artifact.
Models. We evaluate the robustness of VLMs to artifact-based attacks by varying vision-language fusion strategy, vision encoder architecture, and pretraining dataset size and curation. For fusion, we test contrastive learning (CLIP) [26], sigmoid-based alignment (SigLIP) [35], and LLM-based fusion (LLaVA) [18]. To assess the impact of architecture size, we evaluate ViT-B-32, ViT-B-16, ViT-L-14. Finally, we compare models pretrained on LAION-2B [28], and DataComp [10] to examine whether dataset curation affects attack vulnerability.
Artifact Setting. We fix the artifact size to 10th of the image size and transparency to 1.0. Refer to Supplementary. D for a visual example and ablation of both settings.
@@ 4.1 Results
We benchmark the effect of artifact-based attacks on Vision-Language Models (VLMs) across different vision-language fusion methods, model architectures, and pretraining datasets. We make the following observations:
@@
Artifact Attacks are Effective across Diverse Model Configurations. As seen in Fig. 4, artifact-based attacks consistently achieve high success rates across different vision-language fusion strategies, pretraining datasets, and model architectures, reaching up to 90% success. Comparing the artifacts, text-based attacks are the most effective, followed by graphics with embedded text, and finally graphics without text, suggesting that models are most vulnerable to explicit textual cues but still susceptible to purely visual artifacts.
@@
CLIP is most vulnerable to Artifact Attacks. In Fig. 4(a), models trained with standard contrastive learning (OpenAI CLIP) exhibit the highest attack success rates across all artifact types, particularly for text-based artifacts, reinforcing CLIP’s strong reliance on textual cues. The LLM-based fusion model (LLaVA) shows a moderate reduction in vulnerability, likely due to its alignment on a smaller, curated dataset such as COCO [16], which likely have much fewer graphics and logos. Similarly, SigLIP, which employs a different training objective, achieves moderate reduction in vulnerability too on on par with LLaVA. This might be due to SigLIP’s independent binary classification approach which may limit the reinforcement of spurious associations. However, its vulnerability remains comparable to LLaVA, indicating that while altering the training objective helps, it does not fully eliminate susceptibility to artifact-based attacks.
@@
Pretraining Dataset Curation has Minimal Effect on Success Rate. Fig. 4(b) shows that models pretrained DataComp [10] exhibit a slightly improved robusteness to artifact attacks than uncurated Laion400M. Despite DataComp’s filtering, it does not majorly reduce susceptibility, suggesting that its filtering strategy may be ill-suited for mitigating spurious correlations. DataComp primarily optimizes for image quality and semantic similarity to high-quality benchmarks, rather than explicitly removing textual artifacts or graphical elements that reinforce spurious correlations. As a result, while the dataset curation improves image-text alignment for downstream tasks, it fails to eliminate the kind of incidental patterns that make models vulnerable to artifact-based attacks. This highlights a critical limitation: filtering strategies that focus on dataset relevance may not address artifact-based attacks.
@@
Larger Architecture is More Vulnerable to Textual Attacks. As seen in Fig. 4(c), larger models, such as ViT-L-14, exhibit higher attack success rates on text-based attacks. This can be attributed to their increased capacity for optical character recognition, enabling them to more effectively detect and leverage textual elements in images. As a result, they become more susceptible to attacks that exploit text-based spurious correlations present in web-scale datasets.
@@
Figure 5: Confidence in Target Class when different types of artifacts—graphics with text (orange), graphics without text (green), and text (red)—are introduced, compared to the baseline with no artifacts (blue). Performance is averaged across all models and datasets. Results demonstrate that artifacts significantly increase model confidence in the target class.
@@
Figure 6: Success Rate of Combining artifacts. Performance is averaged across all models and datasets. Results indicate that combining artifact types leads to higher attack success.
@@
Figure 7: Sample Size vs Success Rate. Success rate of artifact-based attacks vs the number of samples used from the downstream dataset to estimate their effect. Performance is averaged across all models and datasets. Results indicate that a relatively small sample size is sufficient to assess attack effectiveness.
@@
Artifacts Increase Model Confidence in the Target Class. Fig. 5 shows that the introduction of textual and graphical artifacts significantly increases the model’s confidence in the target class compared to the no-artifact baseline. Text-based artifacts lead to the highest confidence scores, followed by graphics with embedded text and graphics without text. This increased confidence correlates with the attack success rates observed in prior experiments, suggesting that artifacts not only influence predictions but also enhance model certainty in incorrect classifications.
@@
The Compounding Effect of Different types of Artifacts Fig. 6 presents the attack success rate across different artifact types and their combinations. The results highlight a clear synergistic effect when multiple artifacts are used together, significantly increasing misclassification rates compared to individual artifacts. While text-only artifacts remain highly effective, their impact is amplified when combined with graphical symbols. The highest success rates are observed in (A + C) and (A + B) cases, where text is paired with either graphical symbols with or without text, reaching near 100% attack success in some instances. This suggests that models do not treat textual and graphical cues independently but rather reinforce their reliance on both when they co-occur, exacerbating spurious correlations learned during pretraining. Additionally, graphical artifacts without text (B) alone exhibit moderate attack success, but when combined with text (B + C), their effectiveness increases substantially.
@@
Figure 8: Transferability of Artifact-Based Attacks Across Different Vision Encoder Architectures. The heatmap shows the success rate (%) of artifacts search on a specific model (y-axis) and evaluated on another model (x-axis). Higher success rates indicate greater transferability of spurious correlations. Results indicate that artifacts effect is highly transferable across architectures.
@@
Figure 9: Incorporating Artifact Description into the Prompt. Results indicate that the attack effect can be significantly mitigated with more informative prompts.
@@
A Small Sample Size is Enough for Estimating Artifact Effect. A key question for our attacks is how many samples are needed to reliably estimate an artifact’s attack effect. Fig. 7 shows that across all artifact types, the attack success rate stabilizes quickly as the number of samples increases, indicating that only a small sample size is required to obtain a reliable estimate. This result is significant as it suggests that evaluating artifact-based attacks does not require extensive sampling, which greatly reduces computational overhead. By using only a fraction of the dataset, we can efficiently assess the impact of artifacts while preserving computational resources. This finding is particularly important for large-scale evaluations, where repeatedly testing artifacts on full datasets would be prohibitively expensive.
@@
Artifact-Based Attacks are Highly Transferable. Fig. 8 illustrates that artifact-based attacks are highly transferable across different models. Across all artifact types, success rates remain consistently high when applying artifacts mined from one model to another. The most significant drop in success rate occurs when transferring attacks to a model trained with SigLIP [35], which employs a different training objective. However, even in this case, the attack success rates remain substantial, indicating that spurious correlations exploited by these attacks persist across diverse model architectures and training paradigms. This suggests that mitigating these vulnerabilities will require more than simply altering the training objective.
5 Mitigating Artifact Attacks by Incorporating Artifact Descriptions
@@
Recent work [7] has uncovered an important insight: VLMs and LVLMs can be guided to adjust their attention toward specific visual elements when provided with more informative prompts. This finding was leveraged as a mitigation strategy for typographic attacks by explicitly incorporating the attack text into the prompt. For example, assume the an image of a dog is attacked with the text of "a cat", then rather than using a "a photo of a dog", the authors use: "a photo of a dog with cat written on it". While this approach is effective for non-matching text artifacts, it does not generalize to graphical artifacts, which lack an explicit textual form and cannot be directly incorporated into the language input.
To address this limitation, we propose using structured descriptions of graphical symbols to make these artifacts explicit to the model. Specifically, we introduce an artifact-aware description augmentation approach that employs a dedicated detection pipeline: OCR extracts textual elements, while a separate detection model identifies logos and other graphical symbols. The extracted artifacts are then transformed into descriptions and appended to the input, ensuring that the model processes them explicitly rather than forming unintended associations. Refer to Supplementary E for an example.
Following [7], we leverage the generative capabilities for LVLMs like LLaVA [18] to integrate artifact awareness directly. Instead of relying on an external detection pipeline, we prompt LLaVA to describe any graphical or textual artifacts present in the image, ensuring that these features are explicitly considered during inference.
Results. Fig. 9 shows the effectiveness of our approach by comparing two prompting strategies: one without artifact-related information (Prompt wo/ Artifact Info) and one that explicitly includes artifact details (Prompt w/ Artifact Info). Across all artifact types, artifact-aware prompting successfully reduces attack success rates, demonstrating that explicitly instructing the model to account for artifacts can weaken their influence. Despite this improvement, text-based artifacts remain the most effective, even under artifact-aware prompting, reinforcing the strong textual bias in vision-language models. Graphics with embedded text also show a significant drop in success rate, indicating that their impact is closely tied to textual cues. However, graphical artifacts without text exhibit the least reduction in attack success rate, suggesting that these artifacts are more insidious, as they do not rely on direct textual manipulations and remain a greater threat. These results highlight the need for more robust defenses beyond simple prompt modifications, as visual artifacts can still persistently mislead model predictions.
@@ 6 Conclusion
In this work, we introduced Artifact-based Attacks, a novel vector of attacks that make use of non-matching text and graphical artifacts—both with and without embedded text— to mislead model predictions, revealing vulnerabilities beyond traditional typographic attacks. Our results show that these attacks not only increase model confidence in incorrect classifications but also persist across different model architectures, vision-language fusion strategies, and pretraining datasets. Moreover, we showed how our search process for attack artifacts is efficient only requiring a handful of samples to approximate attack effect. While mitigation strategies such as artifact-aware prompting reduce attack success rates, they are insufficient to eliminate the threat, particularly for non-textual graphical artifacts.
@@ Appendix A More examples of Attack Artifacts
In Sec. 3.2, we discuss insights from the found attack artifacts. In this Section, we offer more examples and discussion. Fig. 10 reveals consistent patterns in the types of artifacts that successfully manipulate model predictions. Text-based artifacts frequently include words that share phonetic, visual, or partial textual similarities with the target class. For instance, "KID" appears as an artifact for Child, likely due to its strong semantic association, while "BING" is linked to Boeing, exploiting visual resemblance to the company name. Similarly, artifacts like "K-POP" for South Korea and "Small" for Smiling demonstrate how models latch onto common co-occurring words rather than genuine visual cues.
Graphics with embedded text also play a significant role in misleading models. These artifacts often feature brand names, stylized logos, or generic text associated with the target category. For instance, the inclusion of "Abacus" and "Arbys" as artifacts for Airbus suggests that the model has learned to associate certain brand names or typographic styles with aircraft manufacturers. Likewise, "Happy Bakery Cafe" and "Smile!" appearing under Smiling indicate that models are influenced by commercial logos and positive branding elements rather than facial features.
@@
Across all target classes, these results highlight that artifact-based attacks exploit both direct textual similarities and broader visual associations, making them an effective and adaptable attack vector. The presence of corporate logos, branding, and culturally specific symbols (e.g., the Brazilian flag colors, K-Pop branding for South Korea) suggests that models are influenced by common internet-scale data distributions rather than purely semantic understanding. This demonstrates the pervasive reliance on spurious correlations, emphasizing the need for more robust dataset curation and training strategies to mitigate these vulnerabilities.

### v2

@@ 3.3 Attack Success Evaluation
Given an input image $x$ and an artifact $a$ placed at location $l$, we define the attack-modified image as $x^{a}=x\oplus a$, where $\oplus$ denotes artifact insertion. To verify that misclassification stems from the artifact itself, we introduce a masking-based validation step. We generate a masked version of the image, $x^{\neg l}$, where the region $l$ is occluded using a mask $M_{l}$.
@@
Figure 4: Breakdown of Artifact Attacks Performance. Our Web Artifact attacks are highly effective across different configurations: namely (a) Vision/Language Fusion, (b) Pretraining Dataset, and (c) Vision Encoder Architecture. Refer to Sec.˜4.1 for further discussion..
@@ 4 Experiments
Datasets. We use CC12M [6] as our primary dataset for searching and extracting artifacts. Our extraction process yields 87k artifacts, but this approach is dataset-agnostic and can be applied to any large-scale image-text dataset, such as LAION [30, 31]. To evaluate the effectiveness of these artifacts, we conduct experiments across five distinct tasks spanning both human-related attributes and object/geography classification. For human-related tasks, we predict Gender (Male/Female) and Age (Child, Teen, Adult, Elderly) using FairFace [15], as well as Smiling using CelebA [20]. For non-human classification tasks, we predict aircraft models using FGVC Aircraft [21] and countries using Country211 [28]. For each downstream dataset, we use only 32 images per class to estimate the Artifact Success Rate (ASR), which as we demonstrate in Sec.˜4.1, is sufficient to estimate ASR.
Metrics. We use the success rate as outlined in Sec.˜3.3. In short, given a target class $c^{*}$, we measure attack effectiveness by evaluating whether the model misclassifies an image after the artifact that correlates with $c^{*}$ is introduced while ensuring that the image remains correctly classified when the artifact is masked. We also report the change of the confidence score in target class $c^{*}$ after pasting the artifact.
Models. We evaluate the robustness of VLMs to Web Artifact Attacks by varying vision-language fusion strategy, vision encoder architecture, and pretraining dataset size and curation. For fusion, we test contrastive learning (CLIP) [28], sigmoid-based alignment (SigLIP) [37], and LLM-based fusion (LLaVA) [19]. To assess the impact of architecture size, we evaluate ViT-B-32, ViT-B-16, ViT-L-14 while we fix the pretraining dataset to LAION 2B [30]. Finally, we compare models pretrained on LAION 400M, LAION-2B [30], and DataComp [11] while fixing the pretraining architecture to ViT-B-32 and ViT-B-16 to examine whether dataset curation affects attack vulnerability.
Artifact Setting. We fix the artifact size to 10th of the image size. Refer to Supplementary. D for a visual example and ablation of artifact size and transparency.
@@ 4.1 Results
We benchmark the effect of Web Artifact Attacks on Vision-Language Models (VLMs) across different vision-language fusion methods, model architectures, and pretraining datasets. We make the following observations:
@@
Artifact Attacks are Effective across Diverse Model Configurations. As seen in Fig.˜4, Web Artifact Attacks consistently achieve high success rates across different vision-language fusion strategies, pretraining datasets, and model architectures, reaching up to 90% success. Comparing the artifacts, text-based attacks are the most effective, followed by graphics with embedded text, and finally graphics without text, suggesting that models are most vulnerable to explicit textual cues but still susceptible to purely visual artifacts.
@@
CLIP is most vulnerable to Artifact Attacks. In Fig.˜4(a), models trained with standard contrastive learning (OpenAI CLIP) exhibit the highest attack success rates across all artifact types, particularly for text-based artifacts, reinforcing CLIP’s strong reliance on textual cues. The LLM-based fusion model (LLaVA) shows a moderate reduction in vulnerability, likely due to its alignment on a smaller, curated dataset such as COCO [17], which likely have much fewer graphics and logos. Similarly, SigLIP, which employs a different training objective and a pretraining dataset, achieves moderate reduction in vulnerability too on par with LLaVA.
@@
Pretraining Dataset Curation has Minimal Effect on Success Rate. Fig.˜4(b) shows that models pretrained DataComp [11] exhibit a slightly improved robusteness to artifact attacks than uncurated Laion400M and LAION 2B. Despite DataComp’s filtering, it does not majorly reduce susceptibility, suggesting that its filtering strategy may be ill-suited for mitigating spurious correlations. DataComp primarily optimizes for image quality and semantic similarity to high-quality benchmarks, rather than explicitly removing textual artifacts or graphical elements that reinforce spurious correlations. As a result, while the dataset curation improves image-text alignment for downstream tasks, it fails to eliminate the kind of incidental patterns that make models vulnerable to Web Artifact Attacks. This highlights a critical limitation: filtering strategies that focus on dataset relevance may not address Web Artifact Attacks.
@@
Larger Architecture is More Vulnerable to Textual Attacks. As seen in Fig.˜4(c), larger models, such as ViT-L-14, exhibit higher attack success rates on text-based attacks. This can be attributed to their increased capacity for optical character recognition, enabling them to more effectively detect and leverage textual elements in images. As a result, they become more susceptible to attacks that exploit text-based spurious correlations present in web-scale datasets.
Figure 5: Comparison to Prior Work Attacks. Success rate of Web Artifact attacks vs prior work Typographic Attacks [3].
@@
Figure 6: Confidence in Target Class when different types of artifacts—graphics with text (orange), graphics without text (green), and text (red)—are introduced, compared to the baseline with no artifacts (blue). Results demonstrate that artifacts significantly increase model confidence in the target class.
@@
Figure 7: Success Rate of Combining artifacts. Results indicate that combining artifact types leads to higher attack success.
@@
Figure 8: Sample Size vs Success Rate. Success rate of artifact-based attacks vs the number of samples used from the downstream dataset to estimate their effect. Results indicate that a relatively small sample size is sufficient to assess attack effectiveness.
Artifacts Based Attacks vs Prior Work’s Typographic Attacks. Fig.˜5 presents the average performance of our web-scale Web Artifact Attacks compared to Typographic Attacks [3]. The results demonstrate that our text-based attacks, derived from mined logo datasets, are more effective than typographic attacks, which are typically limited to using text from the opposing class. This suggests that optimizing attacks beyond simple class-matching text—by leveraging a broader range of textual configurations—can significantly enhance their effectiveness. In contrast, graphical attacks tend to be less effective overall. However, they open up a novel attack surface that, as discussed in Section 5, proves more challenging to defend against.
@@
Artifacts Increase Model Confidence in the Target Class. Fig.˜6 shows that the introduction of textual and graphical artifacts significantly increases the model’s confidence in the target class compared to the no-artifact baseline. Text-based artifacts lead to the highest confidence scores, followed by graphics with embedded text and graphics without text. This increased confidence correlates with the attack success rates observed in prior experiments, suggesting that artifacts not only influence predictions but also enhance model certainty in incorrect classifications.
@@
The Compounding Effect of Different Types of Artifacts Fig.˜7 presents the attack success rate across different artifact types and their combinations. The results highlight a clear synergistic effect when multiple artifacts are used together, significantly increasing misclassification rates compared to individual artifacts. While text-only artifacts remain highly effective, their impact is amplified when combined with graphical symbols. The highest success rates are observed in (A + C) and (A + B) cases, where text is paired with either graphical symbols with or without text, reaching near 100% attack success in some instances. This suggests that models do not treat textual and graphical cues independently but rather reinforce their reliance on both when they co-occur, exacerbating spurious correlations learned during pretraining. Additionally, graphical artifacts without text (B) alone exhibit moderate attack success, but when combined with text (B + C), their effectiveness increases substantially.
@@
Figure 9: Transferability of Web Artifact Attacks Across Different Vision Encoder Architectures. The heatmap shows the success rate (%) of artifacts search on a specific model (y-axis) and evaluated on another model (x-axis). Higher success rates indicate greater transferability of spurious correlations. Results indicate that artifacts effect is highly transferable across architectures.
@@
Figure 10: Artifact Aware Prompting. Results indicate that the attack effect can be mitigated with more informative prompts.
@@
A Small Sample Size is Enough for Estimating Artifact Effect. A key question for our attacks is how many samples are needed per class to reliably estimate an artifact’s attack effect. Fig.˜8 shows that across all artifact types, the attack success rate stabilizes quickly as the number of samples increases, indicating that only a small sample size is required to obtain a reliable estimate. This result is significant as it suggests that evaluating Web Artifact Attacks does not require extensive sampling, which greatly reduces computational overhead. By using only a fraction of the dataset, we can efficiently assess the impact of artifacts while preserving computational resources. This finding is particularly important for large-scale evaluations, where repeatedly testing artifacts on full datasets would be prohibitively expensive.
@@
Web Artifact Attacks are Highly Transferable. Fig.˜9 illustrates that Web Artifact Attacks are highly transferable across different models. Across all artifact types, success rates remain consistently high when applying artifacts mined from one model to another. The most significant drop in success rate occurs when transferring attacks to a model trained with SigLIP [37], which employs a different training objective. However, even in this case, the attack success rates remain substantial, indicating that spurious correlations exploited by these attacks persist across diverse model architectures and training paradigms. This suggests that mitigating these vulnerabilities will require more than simply altering the training objective.
5 Mitigation w/ Artifact Aware Prompting
@@
Recent work [8] has uncovered an important insight: VLMs can be guided to adjust their attention toward specific visual elements when provided with more informative prompts. This finding was leveraged as a mitigation strategy for typographic attacks by explicitly incorporating the attack text into the prompt. For example, assume the an image of a dog is attacked with the text of “a cat,” then rather than using “a photo of a dog,” the authors use: “a photo of a dog with cat written on it.” While this approach is effective for textual artifacts, it does not generalize to graphical artifacts, which lack an explicit textual form and cannot be directly incorporated into the language input.
To address this limitation, we propose a mitigation approach inspired by  [8] for the two types of models. For Contrastive VLM(s): we use structured descriptions of graphical symbols to make these artifacts explicit to the model. Specifically, we employ a captioning model to transform the graphics into a textual description which then are appended to the input prompt, ensuring that the model processes them explicitly. Refer to Supp. E for an example. For Large VLM(s): following [8], we leverage the generative capabilities for LVLMs like LLaVA [19] to integrate artifact awareness directly. We prompt LLaVA to describe any graphical or textual artifacts present in the image, ensuring that these features are explicitly considered during inference. Note that our approach does not require model retraining or architectural modifications. The only overhead is the cost of executing the captioning model.
Results. Fig.˜10 illustrates the effectiveness of our approach in reducing attack success rates. Across all artifact types, artifact-aware prompting consistently lowers the success rate, indicating that explicitly guiding the model to consider artifacts can diminish their impact. However, this defense is less effective against our attacks compared to prior typographic attacks [3], with the largest performance gap observed for graphical attacks. This highlights the increased difficulty posed by our attacks and suggests promising directions for future research in developing more robust defenses.
@@ 6 Conclusion
In this work, we introduced Web Artifact Attacks, a novel vector of attacks that make use of non-matching text and graphical artifacts—both with and without embedded text— to mislead model predictions, revealing vulnerabilities beyond traditional typographic attacks. Our results show that these attacks not only increase model confidence in incorrect classifications but also persist across different model architectures, vision-language fusion strategies, and pretraining datasets. Moreover, we showed how our search process for attack artifacts is efficient only requiring a handful of samples to approximate attack effect. While mitigation strategies such as artifact-aware prompting reduce attack success rates, they are insufficient to eliminate the threat, particularly for non-textual graphical artifacts.
@@
Acknowledgments This material is based upon work supported, in part, by DARPA under agreement number HR00112020054. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the supporting agencies.
@@ Appendix A More examples of Attack Artifacts
In Sec. 3.2, we discuss insights from the found attack artifacts. In this section, we offer more examples and discussion. Fig.˜11 reveals consistent patterns in the types of artifacts that successfully manipulate model predictions. Text-based artifacts frequently include words that share phonetic, visual, or partial textual similarities with the target class. For instance, “KID” appears as an artifact for Child, likely due to its strong semantic association, while “BING” is linked to Boeing, exploiting visual resemblance to the company name. Similarly, artifacts like “K-POP” for South Korea and “Small” for Smiling demonstrate how models latch onto common co-occurring words rather than genuine visual cues.
Graphics with embedded text also play a significant role in misleading models. These artifacts often feature brand names, stylized logos, or generic text associated with the target category. For instance, the inclusion of “Abacus” and “Arbys” as artifacts for Airbus suggests that the model has learned to associate certain brand names or typographic styles with aircraft manufacturers. Likewise, “Happy Bakery Cafe” and “Smile!” appearing under Smiling indicate that models are influenced by commercial logos and positive branding elements rather than facial features.
@@
Across all target classes, these results highlight that Web Artifact Attacks exploit both direct textual similarities and broader visual associations, making them an effective and adaptable attack vector. The presence of corporate logos, branding, and culturally specific symbols (e.g., the Brazilian flag colors, K-Pop branding for South Korea) suggests that models are influenced by common internet-scale data distributions rather than purely semantic understanding. This demonstrates the pervasive reliance on spurious correlations, emphasizing the need for more robust dataset curation and training strategies to mitigate these vulnerabilities.